Solve the following equations:
3s = 0
Solve the following equations:
3s = 0

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

We have to divide both sides of the equation by 3,
Then, we get,
= 3s/3 = 0/3
= s= 0

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We know that any number multiplied or divided by 0 gives 0.
So, 
3S = 0
S = 0/3 
S = 0
 
Hope it helps!
Thank you ✌️